# On the Trade-Off between Stability and Fidelity of Gaussian-Smoothed Saliency Maps

### v1

@@ Abstract
Gradient-based saliency maps have been widely used to interpret the decisions of neural network classifiers and discover phenomena from their learned functions. Standard gradient-based maps are frequently observed to be highly sensitive to the randomness of training data and the stochasticity in the training process. In this work, we study the role of randomized smoothing in the well-known Smooth-Grad algorithm in the stability of the gradient-based maps to the randomness of training samples. We extend the algorithmic stability framework to gradient-based saliency maps and prove bounds on the stability error of standard Simple-Grad, Integrated-Gradients, and Smooth-Grad saliency maps. Our theoretical results suggest the role of Gaussian smoothing in boosting the stability of gradient-based maps to the randomness of training settings. On the other hand, we analyze the faithfulness of the Smooth-Grad maps to the original Simple-Grad and show the lower fidelity under a more intense Gaussian smoothing. We support our theoretical results by performing several numerical experiments on standard image datasets. Our empirical results confirm our hypothesis on the fidelity-stability trade-off in the application of Gaussian smoothing to gradient-based interpretation maps.
@@ 1 Introduction
Figure 1: The stability-fidelity trade-off introduced by Gaussian smoothing. We trained neural networks on two disjoint training set splits of ImageNet, and computed the Simple-Grad and Smooth-Grad maps for the same test sample. The similarity between the two maps was higher in Smooth-Grad, while Smooth-Grad loses faithfulness to the Simple-Grad map.
@@
Deep learning models have attained state-of-the-art results over a wide array of supervised learning tasks including image classification [19], speech recognition [12], and text categorization [25]. The trained deep neural networks have been utilized not only to address the target classification task but also to discover the underlying rules influencing the label assignment to a feature vector. To this end, saliency maps, which highlight the features influencing the neural network’s predictions, have been widely used to gain an understanding of phenomena from data-driven models [9] and further to find new discoveries and insights. For example, [26] shows the application of saliency maps to identify the region in fundus images related to anemia, and [4] demonstrates that saliency maps can assist clinicians in diagnosing knee injury from MRI scans.
@@
Specifically, gradient-based maps have been widely applied to compute saliency maps for neural net classifiers. Standard gradient-based saliency maps such as Simple-Grad [30] and Integrated- Gradients [33], represent the input-based gradient of a neural network at a test sample, which reveals the input features with a higher local impact on the neural network classifier’s output. Therefore, the assignment of feature importance scores by a gradient-based saliency map can be utilized to reveal the main features influencing the phenomenon.
@@
However, a consideration for drawing conclusions from gradient-based maps is their potential sensitivity to the training algorithm of the neural network classifier. [2, 35] have shown that gradient maps could be significantly different for independent yet identically distributed training sets, or for a different random initialization used for training the classifier. The sensitivity of the saliency maps is indeed valuable in debugging applications, where the map is used to identify potential reasons behind a misclassification case. On the other hand, the instability of gradient-based maps could hinder the applications of Simple-Grad and Integrated-Gradients maps for phenomena discovery [2], as for a generalizable understanding of the underlying phenomenon, the resulting saliency maps need to have limited dependence on the stochasticity of training data sampled from the underlying data distribution.
@@
In this work, we specifically study the influence of Gaussian smoothing used in the Smooth-Grad algorithm [31] on the stability of a gradient-based map. We provide theoretical and numerical evidence that Gaussian smoothing could increase the stability of saliency maps to the stochasticity of the training setting, which can improve the reliability of applying the gradient-based map for phenomena understanding using the neural net classifier.
@@
To theoretically analyze the stability properties of saliency maps, we utilize the algorithmic stability framework in [5] to quantify the expected robustness of gradient maps to the stochasticity of a randomized training algorithm, e.g. stochastic gradient descent (SGD), and randomness of training data drawn from the underlying distribution. Following the analysis in [13], we define the stability error of a stochastic training algorithm and bound the error in terms of the number of training data and SGD training iterations for the Simple-Grad, Integrated-Gradients, and Smooth-Grad. Our stability error bounds suggest the improvement in stability by applying the randomized smoothing mechanism in Smooth-Grad. Specifically, our error bounds improve by a factor $1/\sigma$ under a standard deviation parameter $\sigma$ of the Smooth-Grad’s Gaussian noise.
@@
In addition to the stability of gradient maps under Gaussian smoothing, we also bound the faithfulness of the Gaussian smoothed saliency maps to the original Simple-Grad and Integrated-Gradients interpretation maps. We show that by increasing the standard deviation parameter $\sigma$, the fidelity error grows proportionally to parameter $\sigma$. This relationship highlights a trade-off between the stability and fidelity of saliency maps, indicating the stabilization offered by Gaussian smoothing at the price of a higher fidelity error compared to the original saliency map. As illustrated in Figure 1, the Smooth-Grad saliency maps could be considerably more stable to the training set of the neural network, while they could signifinctly differ from the original Simple-Grad map. We note that different applications of saliency maps may prioritize stability or fidelity differently. For instance, debugging misclassifications would require higher fidelity, while phenomena discovery would prioritize higher stability. Therefore, understanding the stability-fidelity properties of Smooth-Grad maps helps with a principled application of the algorithm to different tasks.
@@ 2 Related work
Gradient-based Interpretation: A prevalent method for generating saliency maps involves computing the gradient of a deep neural network’s output with respect to an input image. This technique is extensively utilized in numerous related studies, including Smooth-Grad [31], Integrated-Gradients [33], DeepLIFT [29], Grad-CAM [28], and Grad-CAM++ [6]. However, gradient-based saliency maps often exhibit considerable noise. To address this, various methods have been proposed to enhance the quality of these maps by reducing noise. Common strategies include altering the gradient flow through activation functions [36, 32], eliminating negative or minor activations [32, 16, 15], and integrating sparsity priors [20, 37]. Despite these advancements, these techniques do not explicitly tackle the noise and variability introduced by stochastic optimization and the randomness of training data.
@@
Stability Analysis for Interpretation Maps: Beyond visual quality, several studies have explored the stability of interpretation methods. [2] examined the consistency within the same architecture and across different architectures for various interpretation methods on medical imaging datasets, discovering that most methods failed in their tests. [35] conducted experiments revealing that neural networks with different initializations produce distinct saliency maps. Similarly, [8] highlighted the instability in interpretation maps and proposed a metric to evaluate the generalizability of these maps. To address this instability, [35] suggested a Bayesian marginalization technique to eliminate noise from random initialization and stochastic training, though it is computationally intensive due to the need to train multiple networks. Conversely, our Gaussian smoothing approach enhances stability against randomness from both training data and training processes while being computationally efficient, as it requires only a single trained model.
@@
Sanity checks for saliency maps. Evaluating the application of saliency maps has been studied in several related works. The related work [1] proposes sanity checks for interpretation maps, where the saliency map is supposed to depend on the characteristics of input data and how the label $y$ is determined by the input feature vector $\mathbf{x}$. We note that the algorithmic stability considered in our paper is orthogonal to the sanity checks designed in [1], as the algorithmic stability analysis is performed while leaving the distribution $p_{y|\mathbf{x}}$ of training data unchanged. Therefore, our stability bounds for Smooth-Grad do not imply their success or failure in the sanity checks of [1]. Similarly, the algorithmic stability notion in our work is independent of the robustness of interpretation maps to adversarial perturbations studied by [10] and the insensitivity of the interpretation maps to unrelated features discussed in [17], as our defined algorithmic stability implies neither robustness of the map to adversarially-designed perturbations nor its insensitivity to semantically unassociated features.
@@ 3.1 Supervised Learning and Neural Network Optimization
To train the neural network’s parameters, we consider the standard stochastic gradient descent (SGD) optimizer which performs $T$ iterations of uniformly selecting a training data point $(x_{i},y_{i})$ and using the following rule:
@@
$\mathcal{W}_{t+1}=\mathcal{W}_{t}-\alpha_{t}\nabla_{\mathcal{W}}\widetilde{\ell}(\mathcal{W}_{t},x_{i},y_{i})\;\text{where }\widetilde{\ell}(\mathcal{W},x,y):=\mathbb{E}_{\mathcal{V}\in N(0,\kappa^{2}I)}[\ell(\mathcal{W}+\mathcal{V},x,y)]$ (3)
@@ 3.3 Saliency Maps
Simple-Grad. Simple-Grad [3, 30] calculates the gradient of the logit concerning each input pixel:
@@
Smooth-Grad. Smooth-Grad [31] calculates the mean of the Simple-Grad map evaluated at a perturbed input data with a Gaussian noise
$\operatorname{SmoothGrad}_{A(S)}(x,y)=\mathbb{E}_{z\sim N(0,\sigma^{2}I)}\bigl{[}\nabla_{x}(f_{A(S)}(x+z))_{y}\bigr{]}$ (5)
Integrated-Gradients. Integrated-Gradients [33] calculates the integral of the scaled gradient map from a reference point $x_{0}$ to the given data point $x$:
$\operatorname{IntegratedGrad}_{A(S)}(x,y)=(x-x_{0})\odot\int_{0}^{1}\nabla_{x}(f_{A(S)}(x_{0}+\alpha(x-x_{0})))_{y}\mathrm{d}\alpha.$ (6)
@@ 4 Algorithmic Stability of Saliency Maps
$\epsilon_{\mathrm{stability}}(\operatorname{Sal}):=\sup_{S,S^{\prime},x,y}\mathbb{E}_{A}\Bigl{[}$ $\ell^{\prime}(A(S),x,y)-\ell^{\prime}(A(S^{\prime}),x,y)\Bigr{]}$ (10)
We recall the theorem stating that stability implies generalization in expectation [5]. In our specific setting, we extend the original statement and demonstrate that the generalization error of a saliency map is bounded by the stability error of the saliency map algorithm. This indicates the importance of providing theoretical guarantees for stability error.
@@ 4.1 Stability of Simple-Grad
Similar to [13], we consider using the stochastic gradient descent (SGD) algorithm (can be vanilla or noisy version) and need some assumptions for the training loss function.
@@
The following theorem upper bounds the stability error of Simple-Grad in terms of parameters of the SGD algorithm ($T$ and $c$), the Lipschitz and smoothness constants of the loss function ($L$ and $\beta$), and the quantity related to the neural network ($\Gamma$). In the statements, we refer to both the vanilla and noisy versions as the SGD algorithm.
@@
Suppose Asssumption 3(a) and 3(b) hold. If we run SGD for $T$ steps with a decaying step size $\alpha_{t}\leq c/t$ in iteration $t$, by defining $\Gamma=(2\Delta_{k,0}\sum_{i=1}^{k-1}\Delta_{i,1}C)^{1/(\beta c+1)}(2\Delta_{k,0})^{\beta c/(\beta c+1)}$, we can bound the stability error of Simple-Grad by
$\epsilon_{\mathrm{stability}}(\operatorname{SimpleGrad})$ $\leq\operatorname{Up}(\operatorname{SimpleGrad}):=\frac{1+\beta c}{n-1}(2cL)^{\frac{1}{\beta c+1}}T^{\frac{\beta c}{\beta c+1}}\Gamma$
@@ 4.2 Stability of Smooth-Grad
$\epsilon_{\mathrm{stability}}(\operatorname{SmoothGrad})$ $\leq\operatorname{Up}(\operatorname{SimpleGrad})\times\left(\frac{\Delta_{k,1}}{2\sigma\Delta_{k,0}\sum_{i=1}^{k-1}\Delta_{i,1}}\right)^{\frac{1}{\beta c+1}}$
@@
Comparing Theorem 5 and Theorem 6, we observe that Smooth-Grad exhibits a lower stability error by multiplying a dimension-free constant smaller than 1, which keeps improving as the smoothing factor $\sigma$ increases. Theorem 6 also provides a non-asymptotic dimension-free guarantee that as the $\sigma$ value becomes larger, the stability error vanishes to 00. This suggests that the smoothing filter enhances the stability of the saliency map algorithm.

### v2

# Gaussian Smoothing in Saliency Maps: The Stability-Fidelity Trade-Off in Neural Network Interpretability

@@ Abstract
Saliency maps have been widely used to interpret the decisions of neural network classifiers and discover phenomena from their learned functions. However, standard gradient-based maps are frequently observed to be highly sensitive to the randomness of training data and the stochasticity in the training process. In this work, we study the role of Gaussian smoothing in the well-known Smooth-Grad algorithm in the stability of the gradient-based maps to the randomness of training samples. We extend the algorithmic stability framework to gradient-based interpretation maps and prove bounds on the stability error of standard Simple-Grad, Integrated-Gradients, and Smooth-Grad saliency maps. Our theoretical results suggest the role of Gaussian smoothing in boosting the stability of gradient-based maps to the randomness of training settings. On the other hand, we analyze the faithfulness of the Smooth-Grad maps to the original Simple-Grad and show the lower fidelity under a more intense Gaussian smoothing. We support our theoretical results by performing several numerical experiments on standard image datasets. Our empirical results confirm our hypothesis on the fidelity-stability trade-off in the application of Gaussian smoothing to gradient-based interpretation maps.
@@ 1 INTRODUCTION
Figure 1: The stability-fidelity trade-off introduced by Gaussian smoothing. We trained neural networks on two disjoint training set splits of ImageNet, and computed the Simple-Grad and Smooth-Grad maps for the same test sample.
@@
Deep learning models have attained state-of-the-art results over a wide array of supervised learning tasks including image classification (Krizhevsky et al.,, 2012), speech recognition (Graves et al.,, 2013), and text categorization (Minaee et al.,, 2021). The trained deep neural networks have been utilized not only to address the target classification task but also to discover the underlying rules influencing the label assignment to a feature vector. To this end, saliency maps, which highlight the features influencing the neural network’s predictions, have been widely used to gain an understanding of phenomena from data-driven models (Freiesleben et al.,, 2022) and further to find new discoveries and insights. For example, Mitani et al., (2020) shows the application of saliency maps to identify the region in fundus images related to anemia, and Bien et al., (2018) demonstrates that saliency maps can assist clinicians in diagnosing knee injury from MRI scans.
@@
Specifically, gradient-based maps have been widely applied to compute saliency maps for neural net classifiers. Standard gradient-based saliency maps such as Simple-Grad (Simonyan et al.,, 2013) and Integrated- Gradients (Sundararajan et al.,, 2017), represent the input-based gradient of a neural network at a test sample, which reveals the input features with a higher local impact on the neural network classifier’s output. Therefore, the assignment of feature importance scores by a gradient-based saliency map can be utilized to reveal the main features influencing the phenomenon.
@@
However, a consideration for drawing conclusions from gradient-based maps is their potential sensitivity to the training algorithm of the neural network classifier. Arun et al., (2021); Woerl et al., (2023) have shown that gradient maps could be significantly different for independent yet identically distributed training sets, or for a different random initialization used for training the classifier. The sensitivity of the saliency maps is indeed valuable in debugging applications, where the map is used to identify potential reasons behind a misclassification case. On the other hand, the instability of gradient-based maps could hinder the applications of Simple-Grad and Integrated-Gradients maps for phenomena discovery (Arun et al.,, 2021), as for a generalizable understanding of the underlying phenomenon, the resulting saliency maps need to have limited dependence on the stochasticity of training data sampled from the underlying data distribution.
@@
In this work, we specifically study the influence of Gaussian smoothing used in the Smooth-Grad algorithm (Smilkov et al.,, 2017) on the stability of a gradient-based map. We provide theoretical and numerical evidence that Gaussian smoothing could increase the stability of saliency maps to the stochasticity of the training setting, which can improve the reliability of applying the gradient-based map for phenomena understanding using the neural net classifier.
@@
To theoretically analyze the stability properties of saliency maps, we utilize the algorithmic stability framework in Bousquet and Elisseeff, (2002) to quantify the expected robustness of gradient maps to the stochasticity of a randomized training algorithm, e.g. stochastic gradient descent (SGD), and randomness of training data drawn from the underlying distribution. Following the analysis in Hardt et al., (2016), we define the stability error of a stochastic training algorithm and bound the error in terms of the number of training data and SGD training iterations for the Simple-Grad, Integrated-Gradients, and Smooth-Grad. Our stability error bounds suggest the improvement in stability by applying the randomized smoothing mechanism in Smooth-Grad. Specifically, our error bounds improve by a factor $1/\sigma$ under a standard deviation parameter $\sigma$ of the Smooth-Grad’s Gaussian noise.
@@
In addition to the stability of gradient maps under Gaussian smoothing, we also bound the faithfulness of the Gaussian smoothed saliency maps to the original Simple-Grad and Integrated-Gradients interpretation maps, for which we define fidelity error [^1] In this paper, the fidelity term refers to the faithfulness of the regularized saliency map to the original saliency map.. We show that by increasing the standard deviation parameter $\sigma$, the fidelity error grows proportionally to parameter $\sigma$. This relationship indicates the stabilization offered by Gaussian smoothing at the price of a higher fidelity error compared to the original saliency map. As illustrated in Figure 1, the Smooth-Grad saliency maps could be considerably more stable to the training set of the neural network, while they could significantly differ from the original Simple-Grad map. We note that different applications of saliency maps may prioritize stability or fidelity differently. For instance, debugging misclassifications would require higher fidelity, while phenomena discovery would prioritize higher stability. Therefore, understanding the stability-fidelity properties of Smooth-Grad maps helps with a principled application of the algorithm to different tasks.
@@ 2 RELATED WORK
Gradient-based Interpretation A prevalent method for generating saliency maps involves computing the gradient of a deep neural network’s output with respect to an input image. This technique is extensively utilized in numerous related studies, including Smooth-Grad (Smilkov et al.,, 2017), Integrated-Gradients (Sundararajan et al.,, 2017), DeepLIFT (Shrikumar et al.,, 2017), Grad-CAM (Selvaraju et al.,, 2017), and Grad-CAM++ (Chattopadhay et al.,, 2018). However, gradient-based saliency maps often exhibit considerable noise. To address this, various methods have been proposed to enhance the quality of these maps by reducing noise. Common strategies include altering the gradient flow through activation functions (Zeiler and Fergus,, 2014; Springenberg et al.,, 2014), eliminating negative or minor activations (Springenberg et al.,, 2014; Kim et al.,, 2019; Ismail et al.,, 2021), and integrating sparsity priors (Levine et al.,, 2019; Zhang and Farnia,, 2023). Despite these advancements, these techniques do not explicitly tackle the noise and variability in stochastic optimization and the randomness of training data.
@@
Stability Analysis for Interpretation Maps Beyond visual quality, several studies have explored the stability of interpretation methods. Arun et al., (2021) examined the consistency within the same architecture and across different architectures for various interpretation methods on medical imaging datasets, discovering that most methods failed in their tests. Woerl et al., (2023) conducted experiments revealing that neural networks with different initializations produce distinct saliency maps. Similarly, Fel et al., (2022) highlighted the instability in interpretation maps and proposed a metric to evaluate the generalizability of these maps. To address this instability, Woerl et al., (2023) suggested a Bayesian marginalization technique to eliminate noise from random initialization and stochastic training, though it is computationally intensive due to the need to train multiple networks. Furthermore, Zhang and Farnia, (2023) propose MoreauGrad which is a robust and sparsified version of the SimpleGrad and SmoothGrad algorithms. The related works (Gong et al., 2024a, ; Gong et al.,, 2025) develop adversarial training methods for improving the robustness and visual quality of saliency maps, while Gong et al., 2024b leverage image super-pixels for enhanced stability. On the other hand, our work focuses on the stability effects offered by Gaussian smoothing applied in SmoothGrad.
@@
Sanity checks for saliency maps Evaluating the application of saliency maps has been studied in several related works. The related work (Adebayo et al.,, 2018) proposes sanity checks for interpretation maps, where the saliency map is supposed to depend on the characteristics of input data and how the label $y$ is determined by the input feature vector $\mathbf{x}$. We note that the algorithmic stability considered in our paper is orthogonal to the sanity checks designed in Adebayo et al., (2018), as the algorithmic stability analysis is performed while leaving the distribution $p_{y|\mathbf{x}}$ of training data unchanged. Similarly, the algorithmic stability notion in our work is independent of the robustness of interpretation maps to adversarial perturbations studied by Ghorbani et al., (2019) and the insensitivity of the interpretation maps to unrelated features discussed in Kindermans et al., (2019), as our defined algorithmic stability implies neither robustness of the map to adversarially-designed perturbations nor its insensitivity to semantically unassociated features.
@@ 3.1 Supervised Learning and Neural Network Optimization
To train the neural network’s parameters, we consider the standard stochastic gradient descent (SGD) optimizer that performs $T$ iterations of uniformly selecting a training data point $(x_{i},y_{i})$ and using this rule:
@@
$\mathcal{W}_{t+1}=\,\mathcal{W}_{t}-\alpha_{t}\nabla_{\mathcal{W}}\widetilde{\ell}(\mathcal{W}_{t},x_{i},y_{i})$
where ~$\widetilde{\ell}(\mathcal{W},x,y):=\,\mathbb{E}_{\mathcal{V}\in N(0,\kappa^{2}I)}[\ell(\mathcal{W}+\mathcal{V},x,y)]$ (3)
@@ 3.3 Saliency Maps
Simple-Grad. Simple-Grad (Baehrens et al.,, 2010; Simonyan et al.,, 2013) calculates the gradient of the logit concerning each input pixel:
@@
Smooth-Grad. Smooth-Grad (Smilkov et al.,, 2017) calculates the mean of the Simple-Grad map evaluated at a perturbed input data with a Gaussian noise
$\operatorname{SmoothGrad}_{A(S)}(x,y)$
=\,= $\mathbb{E}_{z\sim N(0,\sigma^{2}I)}\bigl{[}\nabla_{x}(f_{A(S)}(x+z))_{y}\bigr{]}$ (5)
Integrated-Gradients. Integrated-Gradients (Sundararajan et al.,, 2017) calculates the integral of the scaled gradient map from a reference point $x_{0}$ to the given data point $x$:
$\operatorname{IntegratedGrad}_{A(S)}(x,y)$
=\,= $(x-x_{0})\odot\int_{0}^{1}\nabla_{x}(f_{A(S)}(x_{0}+\alpha(x-x_{0})))_{y}\mathrm{d}\alpha.$ (6)
@@ 4 ALGORITHMIC STABILITY OF SALIENCY MAPS
$\epsilon_{\mathrm{stability}}(\operatorname{Sal}):=\sup_{S,S^{\prime},x,y}\mathbb{E}_{A}\Bigl{[}$ $\ell^{\prime}(A(S),x,y)$
$-\ell^{\prime}(A(S^{\prime}),x,y)\Bigr{]}$ (10)
We recall the theorem stating that stability implies generalization in expectation (Bousquet and Elisseeff,, 2002). In our setting, we extend the original statement and demonstrate that the generalization error of a saliency map is bounded by the stability error of the saliency map algorithm. This indicates the importance of providing theoretical guarantees for stability error.
@@ 4.1 Stability of Simple-Grad
Similar to Hardt et al., (2016), we consider using the stochastic gradient descent (SGD) algorithm (can be vanilla or noisy version) and need some assumptions for the training loss function.
@@
The following theorem upper bounds the stability error of Simple-Grad in terms of parameters of the SGD algorithm ($T$ and $c$), the Lipschitz and smoothness constants of the loss function ($L$ and $\beta$), and the quantity related to the neural network ($\Gamma$). Here we refer to both the vanilla and noisy versions as the SGD algorithm.
@@
Suppose Assumption 3(a) and 3(b) hold. If we run SGD for $T$ steps with a decaying step size $\alpha_{t}\leq c/t$ in iteration $t$, by defining $\Gamma=(2\Delta_{k,0}\sum_{i=1}^{k-1}\Delta_{i,1}C)^{1/(\beta c+1)}(2\Delta_{k,0})^{\beta c/(\beta c+1)}$, we can bound the stability error of Simple-Grad by
$\epsilon_{\mathrm{stability}}(\operatorname{SimpleGrad})$ $\leq\operatorname{Up}(\operatorname{SimpleGrad})$
$:=\,\frac{1+\beta c}{n-1}(2cL)^{\frac{1}{\beta c+1}}T^{\frac{\beta c}{\beta c+1}}\Gamma$
@@ 4.2 Stability of Smooth-Grad
$\epsilon_{\mathrm{stability}}(\operatorname{SmoothGrad})$ $\leq\operatorname{Up}(\operatorname{SimpleGrad})$
×$\times\left(\frac{\Delta_{k,1}}{2\sigma\Delta_{k,0}\sum_{i=1}^{k-1}\Delta_{i,1}}\right)^{\frac{1}{\beta c+1}}$
@@
Comparing Theorem 5 and Theorem 6, we observe that Smooth-Grad exhibits a lower stability error by multiplying a dimension-free constant smaller than 1. Theorem 6 also provides a non-asymptotic dimension-free guarantee that as the $\sigma$ value becomes larger, the stability error vanishes to 00. This suggests that the smoothing filter enhances the algorithmic stability of high-dimensional saliency maps.